OpenStudy (anonymous):

Solve 3^x = 1/9 -2 -1/2 1/2

OpenStudy (midhun.madhu1987):

3^x = 3^-2

OpenStudy (anonymous):

so, my answer should be -2?

OpenStudy (akashdeepdeb):

\[3^x = \frac{1}{9}\] \[\frac{1}{9} = \frac{1}{3^2} = 3^{-2}\] \[3^x = 3^{-2}\] What do you think your answer should be? :)

OpenStudy (akashdeepdeb):

@vbryand ?